# First Beam Neutrinos Observed with an LAPPD in the ANNIE Experiment

### v1

@@ 4.2 Selection of Charged-Current Neutrino Interactions
Once an ANNIE LAPPD data frame has been matched to information from the other ANNIE subsystems, information from those subsystems may be used to select the LAPPD data frames most consistent with muons from neutrino interactions either in the tank or in the dirt and rock upstream of the detector (“dirt muons”). The resulting sample of neutrino-induced muons can then be used to study track reconstruction using the LAPPD, and compare it to reconstruction using the other detector subsystems.
The red (striped fill) overlay in Figure 8 shows the sub-sample of matched beam events that contain at least one PMT cluster within the initial 2 s readout window. A PMT cluster represents a set of PMTs that registered light consistent with a single track, shower, or neutron capture gamma cascade. It is defined as a set of at least five PMT hits within 50 ns of one another (50 ns being approximately twice the maximal expected time difference between hits due to light propagation times within the tank). Clusters are found by searching the full readout for the 50 ns window that contains the maximal number of hits. These hits are assigned to a cluster and removed from consideration. The process is then repeated until a full scan of the readout fails to produce five or more hits in any 50 ns window. Selecting only PMT clusters for which the average hit time lies in the initial readout window isolates beam-induced neutrino interactions and eliminates PMT clusters due to later detector activity, such as neutron interactions and muon decays. Here, the term hit is defined based on the techniques for baseline subtraction, pulse finding, and waveform calibration described in Ref [22] for tank PMTs.
@@
where is the integrated charge seen by the ith PMT and is the number of operational PMTs. The for a PMT not in the cluster is treated as zero. This requirement is adapted from the WATCHBOY experiment [23] and suppresses clusters arising from instrumentation effects and noise. The additional impact of the PMT cluster quality cut is illustrated in the blue (cross-hatch) overlay in Figure 8.
Figure 9 shows the further impact of requiring activity in the MRD consistent with a muon arising from a charged-current neutrino interaction. TDC units attached to the MRD scintillator paddles register the time at which a muon has passed through the paddle and produced scintillation light, i.e. an MRD hit. A cluster of at least four MRD hits within 30 ns, two from vertical paddles and two from horizontal paddles, is required to successfully reconstruct a muon track. The muon track reconstruction procedure used for the MRD is discussed in greater detail in Refs [24] and [25]. This cut has the most dramatic impact on the available statistics, since it limits the sample to forward-directed muons energetic enough to penetrate the MRD. In order to select a sample of muons that we also expect to illuminate the LAPPD, we further require that the Cherenkov disk observed by the tank PMTs overlaps with or surrounds the LAPPD (muon topology cut).
@@
Table 1 reports the event yield after each selection cut for the entire sample, and the 1.6 s spill region, as well as the estimated event yield for true beam interactions. The beam event count is determined by requiring events to fall within the spill region, and using a sample well outside the spill region to estimate the background. From this table, we see that the additional background suppression provided by the MRD requirement is a modest ( percentage points in signal purity) improvement over the high-quality PMT cluster requirement alone. Further requiring that the PMT and MRD clusters are within of one another provides another slight boost in signal purity by suppressing random coincidences between beam events and cosmic ray muons. The muon topology cut improves signal purity by an additional percentage point.
These cuts clearly select a high purity sample of muons arising from charged-current neutrino interactions in time with the BNB neutrino beam that illuminate the central ANNIE LAPPD. This not only establishes the first observation of beam neutrinos with an LAPPD, but also shows that we can leverage the full detector capabilities of ANNIE to select a useful calibration sample. With this sample, the response of the LAPPD can be studied in relation to the parameters of a muon neutrino interaction as reconstructed independently by other detector systems.
The selected sample at this point provides a high-purity sample of charged-current neutrino interactions, useful for studying muon reconstruction using the LAPPD. We can select a subset of events from this sample for physics measurements by restricting the sample to neutrino interactions that take place inside the ANNIE detector volume. We do so by requiring that there be no activity in the FMV within the time window ns ns. As expected, this cut has no significant impact on signal purity.
@@ 4.3 A First Look at the Single-Muon LAPPD Response
LAPPDs in ANNIE operate in an intermediate statistical regime. However, we have observed that most visible waveforms do not correspond to a single photoelectron liberated by the photocathode, but are rather a superposition of such pulses. This superposition arises as Cherenkov photons arrive within a time period comparable to the characteristic width of a single-photoelectron pulse in time; moreover, the charge clouds induced by these photoelectrons can spread over multiple microstrips. These features of the data suggest that reconstruction is better done by fitting a holistic model of the voltage-time response rather than through deconvolution and isolation of individual photons. Future work will develop both the fitting method and the necessary modeling of the time-voltage response for different muon parameters in order to benchmark these capabilities.
@@ 5 Conclusions
ANNIE has achieved the first successful deployment and operation of an LAPPD in a running neutrino experiment and detected, for the first time, light from beam neutrino interactions using an LAPPD. With these milestones, ANNIE has moved this technology beyond the domain of testbench studies and detector simulation into the domain of full experimental reality.
Leveraging the additional information from the rest of the ANNIE detector, we have been able to select a 98% pure sample of beam neutrino interactions (1030 events in total) that illuminated the LAPPD. This sample of neutrino events, with accompanying MRD tracks, provides a valuable dataset for understanding and developing track reconstruction using LAPPDs in a real-world physics context. As illustrated by the waveform patterns and timing gradient shown in this paper, an LAPPD produces rich spatio-temporal data with the potential to constrain muon track parameters using a single photosensor. We now have the data required to develop and test the techniques needed to fully exploit that potential. With a full year of exposure and additional LAPPDs, ANNIE is capable of collecting thousands more neutrino interactions with expanded coverage.
Equally important is the knowledge gained from these early deployments about the hardware and design of large LAPPD systems. ANNIE continues to provide the first experience of truly long-term operation of LAPPDs in space-confined conditions and share that experience with both the LAPPD manufacturer and other early adopters of this technology. Our experiences have already suggested improvements to the slow controls, waterproof housing, and electronics readout systems.
ANNIE was built with a major goal of demonstrating ways in which the combination of LAPPDs with chemical enhancements such as Gd and water-based liquid scintillator could fundamentally change how we think about neutrino detection on a water target. With their unique combination of timing and spatial granularity, LAPPDs have the potential to open up next-generation neutrino experiments and collider experiments to expanded physics capabilities, new techniques, and novel detector concepts.
@@ Appendix A Surface-to-Water Cabling
As discussed in Section 3, high-density waterproof cables from Falmat (Falmat XtremeNet) [26] connect the surface rack electronics and power to the PAL. Here we describe in detail the design of this cable system.
@@
The surface end of the waterproof cable terminates the Cat6 twisted-pair data lines in a female RJ-45 jack. A Cat6 cable is used to patch the data lines directly from this jack to the surface components of the LAPPD readout electronics. This maintains the native noise and cross-talk performance of the cables. The power and control lines go to an Amphenol Industrial PT06E-14-12P connector [27] that mates to a connector on the “breakout boxes” that serve power and slow controls at the surface.
The cables are routed to the PAL through rectangular openings (LAPPD deployment ports) at the top of the tank. The PAL end of the Falmat cable is spliced to a waterproof lead cable attached to one half of a floodproof SubConn Power Ethernet connector from MacArtney [28]. The mating half of the SubConn connector is installed in the LAPPD housing. The Falmat XtremeNet cable itself is compatible with Gd-water and is left exposed, but the SubConn connector and lead cable are not as Gd-compatible and have been Teflon-wrapped to prevent leaching.
@@ Appendix B Power Distribution and Slow Controls
As noted in Section 3, the LAPPD subsystem has its own dedicated power supply and distribution system. An Excelsys Xgen XLC[18] power supply system provides the primary power source for the LAPPD subsystem, with two 12V 240W Xg3 modules powering the PALs high- and low-voltage distribution systems. A single 5V 200W Xg2 module powers the ACCs and Raspberry Pis on the surface.
Power is routed and delivered through breakout boxes, which (a) serve power to the underwater PAL modules, (b) receive alarm and monitoring signals from the PALs, and (c) contain two odel B Raspberry Pi 3s [29] that manage slow controls, trigger configuration, and alarms. A single breakout box is designed to service two PALs. In addition to providing an interface for the termination of the waterproof cables and distributing power and communications lines, the Raspberry Pis serve as the interface between the DAQ and LAPPD slow controls. Each Raspberry Pi is outfitted with two add-on mezzanine cards (Pi HATs): a Copperhill technologies PiCAN2 CANBUS interface[30] for communications with the underwater slow controls on the LVHV board and an Electronics-Salon RPi Power Relay Board Expansion Module [31] for shutting off the power to the PAL in the event of a serious water leak. A unit cell of the breakout box and how it maps to the power and slow controls of the PAL is shown in Figure 12.
@@
Control and monitoring of the LVHV board is provided by a Microchip Inc dsPIC33EV256GM106 (PIC) microcontroller[32]. The PIC can control and monitor the low voltage and high voltage circuits on the LVHV, read out onboard humidity and temperature sensors, set thresholds on the trigger board, and communicate with a Raspberry Pi on the surface electronics using a CANBUS interface.
Operation of an ANNIE LAPPD requires a chain of 5 high voltages spanning 3kV-0V. The maximum high voltage for the LAPPD is generated by a single EMCO C40N power converter[33]. This voltage is divided down by a resistive divider network, with parallel zener diodes used to protect against overvoltages across any single stage of the LAPPD. The individual MCPs of the LAPPD are connected with the resistive divider in series. If the individual plate resistances are mismatched, a parallel surface mount resistor is used to equalize the equivalent resistance. Figure 13 shows a schematic of the setup. A single sensor on the LVHV allows read back of a test point that reflects the high voltage value at the top of the stack divided by ten. However, we are unable to interrogate the voltages of each stage independently on the deployed LAPPD in situ.
@@ Acknowledgments
This work was supported by: (1) the U.S. Department of Energy (DOE) Office of Science under award number DE-SC0009999 (UC Davis), award number DE-SC0015684 (Iowa State University), and award number DE-SC0024684 (Florida State University) and by the DOE National Nuclear Security Administration through the Nuclear Science and Security Consortium under award number DE-NA0003180 (UC Davis), (2) Deutsche Forschungsgemeinschaft grants 456139317 (Hamburg) and 490717455 (Mainz and Tübingen), and 552099472 and RTG2796 “Particle Detectors” (Mainz) (3) National Science Foundation Grant No. PHY-2310018 (Ohio State), PHY-2047665 (Rutgers), and OIA-2132223 (South Dakota Mines), (4) UK Research and Innovation FLF MR/S032843/1 (Warwick), (5) Scientific Research Projects (BAP) of Erciyes University under the grant numbers of of FBAU-2023-12325, FDS-2021-11525 and FBA-2022-12207, Work conducted at Lawrence Berkeley National Laboratory performed under the auspices of the U.S. Department of Energy under Contract DE-AC02-05CH11231. Work conducted at Lawrence Livermore National Laboratory was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract DE-AC52-07NA27344, release number LLNL-JRNL-2009446. The work conducted at Brookhaven National Laboratory was supported by the U.S. Department of Energy under contract DE-AC02-98CH10886. The project was funded by the U.S. Department of Energy, National Nuclear Security Administration, Office of Defense Nuclear Nonproliferation Research and Development (DNN R&D). Finally, we gratefully acknowledge all the Fermilab scientists and staff who supported this work through their technical expertise and operational assistance at the Booster Neutrino Beam.

### v2

@@ 4.2 Selection of Charged-Current Neutrino Interactions
In this section, we consider the impact of cuts using information from the entire ANNIE detector that are designed to select charged-current neutrino interactions either in the tank or in the dirt and rock upstream of the detector (“dirt muons”). In evaluating the impact of these cuts on the timing distribution of the matched LAPPD data frames, we see that these cuts preferentially select data frames in time with the spill. An important quality metric for each cut is the sample purity within the beam spill region ( s), that is, the number of events with an LAPPD data frame in the spill region after subtracting the estimated background, divided by the total number of events with an LAPPD data frame in the spill region. We estimate the background rate using two off-beam regions, 1-6 s and 11-20 s. Both the selected spill region and the background estimation regions are shown in Fig. 8.
The red (striped fill) overlay in Figure 8 shows the sub-sample of matched beam events that contain at least one PMT cluster within the initial 2 s readout window. A PMT cluster represents a set of PMTs that registered light consistent with a single track, shower, or neutron capture gamma cascade. It is defined as a set of at least five PMT hits within 50 ns of one another (50 ns being approximately twice the maximal expected time difference between hits due to light propagation times within the tank). Clusters are found by searching the full readout for the 50 ns window that contains the maximal number of hits. These hits are assigned to a cluster and removed from consideration. The process is then repeated until a full scan of the readout fails to produce five or more hits in any 50 ns window. Selecting only PMT clusters for which the average hit time lies in the initial readout window isolates beam-induced neutrino interactions and eliminates PMT clusters due to later detector activity, such as neutron interactions and muon decays. Here, the term hit is defined based on the techniques for baseline subtraction, pulse finding, and waveform calibration described in Ref [Back_2020] for tank PMTs.
@@
where is the integrated charge seen by the ith PMT and is the number of operational PMTs. The for a PMT not in the cluster is treated as zero. This requirement is adapted from the WATCHBOY experiment [DAZELEY2016151] and suppresses clusters arising from instrumentation effects and noise. The additional impact of the PMT cluster quality cut is illustrated in the blue (cross-hatch) overlay in Figure 8.
Figure 9 shows the further impact of requiring activity in the MRD consistent with a muon arising from a charged-current neutrino interaction. TDC units attached to the MRD scintillator paddles register the time at which a muon has passed through the paddle and produced scintillation light, i.e. an MRD hit. A cluster of at least four MRD hits within 30 ns, two from vertical paddles and two from horizontal paddles, is required to successfully reconstruct a muon track. The muon track reconstruction procedure used for the MRD is discussed in greater detail in Refs [Fischer:2020hjj] and [NieslonyThesis]. This cut has the most dramatic impact on the available statistics, since it limits the sample to forward-directed muons energetic enough to penetrate the MRD. In order to select a sample of muons that we also expect to illuminate the LAPPD, we further require that the Cherenkov disk observed by the tank PMTs overlaps with or surrounds the LAPPD (muon topology cut).
@@
Table 1 reports the event yield after each selection cut for the entire sample and for the 1.6 s spill region, as well as the estimated event yield for true beam interactions. From this table, we see that the additional background suppression provided by the MRD requirement is a modest ( percentage points in signal purity) improvement over the high-quality PMT cluster requirement alone. Further requiring that the PMT and MRD clusters are within of one another provides another slight boost in signal purity by suppressing random coincidences between beam events and cosmic ray muons. The muon topology cut improves signal purity by an additional percentage point.
These cuts clearly select a high purity sample of muons arising from charged-current neutrino interactions in time with the BNB neutrino beam that illuminate the central ANNIE LAPPD. This not only establishes the first observation of beam neutrinos with an LAPPD, but also shows that we can leverage the full detector capabilities of ANNIE to select a useful calibration sample. With this sample, the response of the LAPPD can be studied in relation to the parameters of a muon neutrino interaction as reconstructed independently by other detector systems. We can select a subset of events useful for physics measurements by requiring the neutrino interactions take place inside the ANNIE detector volume. We do so by requiring that there be no activity in the FMV within the time window ns ns. As expected, this cut has no significant impact on signal purity.
@@ 4.3 A First Look at the Single-Muon LAPPD Response
We have observed that most visible waveforms in ANNIE LAPPD events do not correspond to a single photoelectron liberated by the photocathode, but are rather a superposition of many such pulses. This superposition arises as Cherenkov photons arrive within a time period comparable to the characteristic width of a single-photoelectron pulse in time; moreover, the charge clouds induced by these photoelectrons can spread over multiple microstrips. These features of the data suggest that reconstruction is better done by fitting a holistic model of the voltage-time response rather than through deconvolution and isolation of individual photons. Future work will develop both the fitting method and the necessary modeling of the time-voltage response for different muon parameters in order to benchmark these capabilities.
[…]
@@ 5 Conclusions
ANNIE has achieved the first successful deployment and sustained operation of a Large Area Picosecond Photodetector (LAPPD) in a running neutrino experiment, and has recorded the first light from beam neutrino interactions with an LAPPD. These results move LAPPDs from test-stand and simulation studies into demonstrated experimental use.
By combining LAPPD data with information from all other ANNIE sub-detector systems, we selected a high-purity sample (98% pure) of beam neutrino interactions (1,045 events in total) that illuminated the LAPPD. This sample, which includes accompanying MRD tracks, provides a realistic testbed for developing and validating LAPPD-based track reconstruction. The waveform morphologies and timing gradients shown here demonstrate that a single LAPPD can deliver rich spatio-temporal information with power to constrain muon track parameters.
The early deployments have also yielded essential system-engineering insights. ANNIE provides the first experience with long-term operation of LAPPDs in a space-constrained water detector and has already informed improvements to slow controls, waterproof housings, and front-end readout electronics. These lessons are directly transferable to future multi-module systems.
With additional LAPPDs and an extended run, ANNIE has collected an order of magnitude more neutrino interactions with greater geometric coverage. That dataset will enable quantitative reconstruction studies and performance benchmarks that are not yet possible with the dataset presented here.
More broadly, ANNIE demonstrates that LAPPDs can change how we approach neutrino detection with water-based targets. The combination of picosecond timing and millimeter-scale spatial sampling opens a path to enhanced physics reach, new reconstruction techniques, and novel detector concepts for next-generation neutrino and collider experiments.
@@ Appendix A Surface-to-Water Cabling
As discussed in Section 3, high-density waterproof cables from Falmat (Falmat XtremeNet) [Falmat] connect the surface rack electronics and power to the PAL. Here we describe in detail the design of this cable system.
@@
The surface end of the waterproof cable terminates the Cat6 twisted-pair data lines in a female RJ-45 jack. A Cat6 cable is used to patch the data lines directly from this jack to the surface components of the LAPPD readout electronics. This maintains the native noise and cross-talk performance of the cables. The power and control lines go to an Amphenol Industrial PT06E-14-12P connector [Amphenol] that mates to a connector on the “breakout boxes” that serve power and slow controls at the surface.
The cables are routed to the PAL through rectangular openings (LAPPD deployment ports) at the top of the tank. The PAL end of the Falmat cable is spliced to a waterproof lead cable attached to one half of a floodproof SubConn Power Ethernet connector from MacArtney [SubConn]. The mating half of the SubConn connector is installed in the LAPPD housing. The Falmat XtremeNet cable itself is compatible with Gd-water and is left exposed, but the SubConn connector and lead cable are not as Gd-compatible and have been Teflon-wrapped to prevent leaching.
@@ Appendix B Power Distribution and Slow Controls
As noted in Section 3, the LAPPD subsystem has its own dedicated power supply and distribution system. An Excelsys Xgen XLC[Excelsys] power supply system provides the primary power source for the LAPPD subsystem, with two 12V 240W Xg3 modules powering the PALs high- and low-voltage distribution systems. A single 5V 200W Xg2 module powers the ACCs and Raspberry Pis on the surface.
Power is routed and delivered through breakout boxes, which (a) serve power to the underwater PAL modules, (b) receive alarm and monitoring signals from the PALs, and (c) contain two model B Raspberry Pi 3s [RPi] that manage slow controls, trigger configuration, and alarms. A single breakout box is designed to service two PALs. In addition to providing an interface for the termination of the waterproof cables and distributing power and communications lines, the Raspberry Pis serve as the interface between the DAQ and LAPPD slow controls. Each Raspberry Pi is outfitted with two add-on mezzanine cards (Pi HATs): a Copperhill technologies PiCAN2 CANBUS interface[Copperhill] for communications with the underwater slow controls on the LVHV board and an Electronics-Salon RPi Power Relay Board Expansion Module [Salon] for shutting off the power to the PAL in the event of a serious water leak. A unit cell of the breakout box and how it maps to the power and slow controls of the PAL is shown in Figure 12.
@@
Control and monitoring of the LVHV board is provided by a Microchip Inc dsPIC33EV256GM106 (PIC) microcontroller[PIC]. The PIC can control and monitor the low voltage and high voltage circuits on the LVHV, read out onboard humidity and temperature sensors, set thresholds on the trigger board, and communicate with a Raspberry Pi on the surface electronics using a CANBUS interface.
Operation of an ANNIE LAPPD requires a chain of 5 high voltages spanning 3kV-0V. The maximum high voltage for the LAPPD is generated by a single EMCO C40N power converter[EMCOC40N]. This voltage is divided down by a resistive divider network, with parallel zener diodes used to protect against overvoltages across any single stage of the LAPPD. The individual MCPs of the LAPPD are connected with the resistive divider in series. If the individual plate resistances are mismatched, a parallel surface mount resistor is used to equalize the equivalent resistance. Figure 13 shows a schematic of the setup. A single sensor on the LVHV allows read back of a test point that reflects the high voltage value at the top of the stack divided by ten. However, we are unable to interrogate the voltages of each stage independently on the deployed LAPPD in situ.
@@ Acknowledgments
This work was supported by: (1) the U.S. Department of Energy (DOE) Office of Science under award number DE-SC0009999 (UC Davis), award number DE-SC0015684 (Iowa State, award number DE-SC0024684 (Florida State University), and DE-SC0014223 (South Dakota School of Mines), and by the DOE National Nuclear Security Administration through the Nuclear Science and Security Consortium under award number DE-NA0003180 (UC Davis), (2) Deutsche Forschungsgemeinschaft grants 456139317 (Hamburg) and 490717455 (Mainz and Tübingen), and 552099472 and RTG2796 “Particle Detectors” (Mainz) (3) National Science Foundation Grant No. PHY-2310018 (Ohio State), PHY-2047665 (Rutgers), and OIA-2132223 (South Dakota Mines), (4) UK Research and Innovation FLF MR/S032843/1 (Warwick), (5) Scientific Research Projects (BAP) of Erciyes University under the grant numbers of FBAU-2023-12325, FDS-2021-11525 and FBA-2022-12207. Work conducted at Lawrence Berkeley National Laboratory was performed under the auspices of the U.S. Department of Energy under Contract DE-AC02-05CH11231. Work conducted at Lawrence Livermore National Laboratory was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under contract DE-AC52-07NA27344, release number LLNL-JRNL-2009446. The work conducted at Brookhaven National Laboratory was supported by the U.S. Department of Energy under contract DE-AC02-98CH10886. The project was supported by the U.S. Department of Energy, National Nuclear Security Administration, Office of Defense Nuclear Nonproliferation Research and Development (DNN R&D). Finally, we gratefully acknowledge all the Fermilab scientists and staff who supported this work through their technical expertise and operational assistance at the Booster Neutrino Beam.